OpenStudy (anonymous):

hey solve for x: sin90x=x^2-2x+2?

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

Remember?: \[\sin 90 = 1\] Therefore x is: \[1 \times x = x^{2} - 2x + 2\] which is equal to: \[x = x ^{2} -2x + 2\] @aryandecoolest : you question's answered

OpenStudy (anonymous):

no no how can you do that 90x itself is an angle.

OpenStudy (mokeira):

sin90x is like \[\sin90\times x\] so u get what im trying to say?

OpenStudy (anonymous):

nopes @Mokeira it's sin(90x)

OpenStudy (mokeira):

exactly it is \[(\sin90)x\]

OpenStudy (mokeira):

ooooooooh. I see what you mean

OpenStudy (anonymous):

no no, don't separate 90 and x.

OpenStudy (mokeira):

Yeah... I got what you are saying..How do you solve it?

OpenStudy (anonymous):

i am confused..you got any idea? @Mokeira

OpenStudy (anonymous):

The thing is if i can get RHS factorized in proper way i can solve by properties sinx=sin(theta)

OpenStudy (mokeira):

 Have you come across something like that?

OpenStudy (anonymous):

naah, can you explain from where it comes?

OpenStudy (mokeira):

It is a cos sin triangle

OpenStudy (mokeira):

I would use that to solve the question

OpenStudy (anonymous):

ok....continue...tell me the final answer.

OpenStudy (anonymous):

actually by hit and trial method i can get answer as 1 like substituting x=1 and then LHS=RHS.

OpenStudy (mokeira):



OpenStudy (mokeira):

i have seen using the Pythagoras theory will not work so forget my triangles

OpenStudy (mokeira):

i dont know how to solve

OpenStudy (anonymous):

well no problem, thanks for your time......proving solves the problem :) answer=1

OpenStudy (mokeira):

i think the sin90x is same as (sin)(90)(X) so you can group as you wish like sin90(x)...then sin90=1 so Sin90x is same as (x)1

OpenStudy (anonymous):

x^2 -2x +2 = (x-1)^2+1 Thus, x^2-2x+2 >=1 we also know, -1<=sin(90x)<=1 so, if sin(90x)=x^2-2x+2 has any solution then it must be sin(90x)=x^2-2x+2=1 now, sin(90x)=1 90x = 90 + 360n , where n is any integer or, x = 1 + 4n also, x^2-2x+2=1 =>x=1 Thus, x=1 is the only solution